OpenStudy (anonymous):

please help ... Write the equation of the line which passes through (2, –1) and is perpendicular to the line with equation 3y – x =

OpenStudy (anonymous):

is this a line graph if it is follow the formula mx + b m standing for gradient, b interception point on the y axis

OpenStudy (anonymous):

3y-x=?

OpenStudy (anonymous):

I guess you're asking for this ? :P ... http://www.okc.cc.ok.us/maustin/EquationsofLine/Equations%20of%20a%20Line.htm here's everything you want to know.

OpenStudy (anonymous):

slope of \(3y-x=\text{whatever}\) is \(\frac{1}{3}\) and so slope of perpendicular line must be the negative reciporcal. the negative reciprocal of \(\frac{1}{3}\) is \(-3\) so you know your slope is \(-3\) and your point is \((2,-1)\) now use the point -slope formula \[y-y_1=m(x-x_1)\] with \[m=-3, x_1=2, y_1=-1\] and you should get your answer

OpenStudy (anonymous):

*negative reciprocal

OpenStudy (anonymous):

you good from there?

OpenStudy (anonymous):

u think so... -1-(-1)=-3(2-2)

OpenStudy (anonymous):

oh actually no, so lets go slow

OpenStudy (anonymous):

your equation should look like either (y=mx+b\) or \(ax+by=c\) one of those two

OpenStudy (anonymous):

in the "point - slope" formula there are three numbers and two variables. the formula is \[y-y_1=m(x-x_1)\] in this formula \(x\) and \(y\) are variables that appear in your answer. the numbers are \(x_1, y_1, m\)

OpenStudy (anonymous):

so when you make the number replacements, leave \(y\) as \(y\) and \(x\) as \(x\) and replace only the \(m\) the \(x_1\) and the \(y_1\) by numbers

OpenStudy (anonymous):

you want to try it or do you want me to write it. your choice

OpenStudy (anonymous):

i want to try it but i am still confused... so x= 2, y= -1

OpenStudy (anonymous):

yes your point is \((2,-1)\) so replace \(x_1\) by 2, replace \(y_1\) by -1, and replace \(m\) by -3 in the formula \(y-y_1=m(x-x_1)\) leave the \(y\) as \(y\) and the \(x\) as \(x\)

OpenStudy (anonymous):

can you wright it out :/

OpenStudy (anonymous):

3x-y=2?

OpenStudy (anonymous):

\[y-y_1=m(x-x_1)\] \[y-(-1)=-3(x-2)\] now there is some algebra to be done. notice that i left the \(x\) and \(y\) exactly where they were, and only replaced the \(m, x_1\) and \(y_1\)

OpenStudy (anonymous):

is it clear how i got that? we knew the three numbers \(m=-3, x_1=2, y_1=-1\) and replaced them in the formula

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok good. now there is some algebra to to before you are done first of all \[y-(-1)=-3(x-2)\] then \[y+1=-3(x-2)\] then distributive property \[y+1=-3x+6\] then subtract 1 to get \[y=-3x+5\]

OpenStudy (anonymous):

if you want it in the other form write it as \[3x+y=5\]

OpenStudy (anonymous):

thank you ! im printing all of these explinations off to help myself and friends